Two jets of material from the center of a radio galaxy fly away in opposite directions. Both

jets move at 0.328 c relative to the galaxy. Determine the magnitude of the velocity of

one jet relative to the other. Correct answer: 0.59228 c.
Answer #1

Relative velocity of one to the other is Ve=u-v luz-V opposite direction 1- UV - u-cu) . uty - 1-uty 1+42 = 24 = 2(0.328c) 14